OpenStudy (anonymous):

What is the area of the circle if XY = 5.4 mm ? A.) 5.4π mm2 B.) 7.29π mm2 C.) 14.58π mm2 D.) 10.8π mm2

OpenStudy (anonymous):

Area Of A Circle=πr^2

OpenStudy (anonymous):

Wait Is There Suppose To Be A Graph?

OpenStudy (anonymous):

no its just a circle look

OpenStudy (anonymous):

Ok, Thats What I Meant, Thank You ☺ So The Formula Is... Area=3.14*radius^2 The Radius Is Half The Circle. XY=The Diameter, You Know What I Mean? So If The Diameter Is 5.4, What Is The Radius?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

Diameter=XY=5.4. What Is The Radius?

OpenStudy (anonymous):

10.

OpenStudy (anonymous):

10.8?

OpenStudy (anonymous):

No, The Diameter Is 5.4, Not The Radius. You Would DIVIDE 5.4 By 2. Not Multiply :P

OpenStudy (anonymous):

2.7

OpenStudy (anonymous):

That Is Correct ☺ Now We Plug It Into The Area Of A Circle Area, \[A= \pi r^2 \] Area=3.14*radius^2 =3.14*(2.7)^2

OpenStudy (anonymous):

Do You Know What To Do Next?

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

Well First You Should Do The Exponents, You Know What An Exponent Is Right?

OpenStudy (anonymous):

so 7.9

OpenStudy (anonymous):

Close, Its 7.29, So Lets Plug That In A=3.14*7.29 Once You Do That, You Have The Area Of A Circle :D

OpenStudy (anonymous):

24.806 is the answer?

OpenStudy (anonymous):

22.8906 Actually, But Just Put 22.89 :P

OpenStudy (anonymous):

but the answer choices are these: A.) 5.4π mm2 B.) 7.29π mm2 C.) 14.58π mm2 D.) 10.8π mm2

OpenStudy (anonymous):

and 22.89 is not there

OpenStudy (anonymous):

I Guess The Answer Is B Then Since \[A=\pi r^2\] Is Pretty Much The Same Thing As \[A=7.29 \pi\] In This Situation

OpenStudy (anonymous):

okay thanks :D

OpenStudy (anonymous):

You're Welcome! Hopefully You Understand This Problem Now :O Good Day ☺

OpenStudy (anonymous):

yes i do :)

OpenStudy (anonymous):

Ok, Good, Tag Me If You Need Anymore Help, Goodbye For Now ☺